# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

### v1

@@ Abstract.
AI code generation tools have gained significant popularity among developers, who use them to assist in software development due to their capability to generate code. Existing studies mainly explored the quality, e.g., correctness and security, of AI-generated code, while in real-world software development, the prerequisite is to distinguish AI-generated code from human-written code, which emphasizes the need to explicitly declare AI-generated code by developers. To this end, this study intends to understand the ways developers use to self-declare AI-generated code and explore the reasons why developers choose to self-declare or not. We conducted a mixed-methods study consisting of two phases. In the first phase, we mined GitHub repositories and collected 613 instances of AI-generated code snippets. In the second phase, we conducted a follow-up industrial survey, which received 111 valid responses. Our research revealed the practices followed by developers to self-declare AI-generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. The reasons for self-declaring AI-generated code include the need to track and monitor the code for future review and debugging, and ethical considerations. The reasons for not self-declaring AI-generated code include extensive modifications to AI-generated code and the developers’ perception that self-declaration is an unnecessary activity. We finally provided guidelines for practitioners to self-declare AI-generated code, addressing ethical and code quality concerns.
@@ 1. Introduction
Artificial Intelligence (AI) techniques, especially those employing Machine Learning (ML) models and Large Language Models (LLMs), are revolutionizing software engineering practices. There has been increased adoption of these models to develop more accurate, faster, and scalable tools that assist software engineers in their development tasks (Yellin, 2023) (e.g., AI-assisted coding (Liang et al., 2024)). One area that has been largely transformed with the advancement in LLMs is automatic code generation (Jiang et al., 2024; Chen et al., 2021). Developers increasingly rely on AI code generation tools in their written code (AIt, 2024; AIR, 2023). The most recent Stack Overflow Developer Survey (2024) shows that there has been a significant adoption of code generation tools in practice (SO-, [n. d.]). This is partially due to the widely available coding-specific tools (e.g., GitHub Copilot (Cop, 2024) and CodeWhisperer (Cod, 2024)) and general-purpose LLMs (e.g., ChatGPT (Cha, 2024) and Gemini (Gem, 2024)) that can generate a large amount of code with little human intervention, faster with high efficiency (Dakhel et al., 2023; Tian et al., 2023).
@@
Figure 1. An example of self-declaration comments of AI-generated code
@@
The main contributions of this work are: (1) a curated dataset of 613 self-declared AI-generated code snippets from open-source repositories on GitHub and performing a comprehensive analysis with a focus on self-declaration comments to understand the practices of self-declaring AI-generated code, (2) an industrial survey to validate the findings of the mining study and explore the potential reasons behind self-declaring (or not) AI-generated code, and (3) guidelines for practitioners to self-declare AI-generated code based on the results of the mining study and the follow-up industrial survey.
Paper Organization: Section 2 discusses related work to this study. Section 3 presents our research questions and outlines the methodology employed for our data collection (from the mining study and the industrial survey). We present our results in Section 4, followed by the discussion of the findings and implications in Section 5. We then present potential threats to the validity in Section 6 followed by the conclusion in Section 7.
@@ 2. Related Work
Quality, Correctness and Security of AI-Generated Code. Several recent studies have also investigated the quality and correctness of AI-generated code. Liu et al. evaluated ChatGPT-generated solutions to programming problems and studied their correctness and quality (through static analysis) (Liu et al., 2024a). The results showed that ChatGPT could generate correct solutions almost 70% of the time but suffers from maintainability issues. Ouyang et al. studied ChatGPT code generation capabilities and found that the non-determinism of ChatGPT affects the correctness and consistency of the generated code (Ouyang et al., 2024). Clark et al. evaluated ChatGPT-generated code quality and found that ChatGPT can generate good quality code through iterative prompts, but the generated code was found to contain bugs (Tambon et al., 2025), citing that ChatGPT-generated code would usually require review before integration into projects. These results show that AI-generated code will not hinder the software development process. Bui et al. introduced a framework (called OPENIA) that uses the internal states of code-generation LLMs to evaluate the correctness of AI-generated code (Bui et al., 2025). This approach was found to outperform existing methods with significant improvements in performance metrics, enabling more effective quality assurance in AI code generation.
@@
Similarly, the security of the generated code is another aspect that has been widely studied. Pearce et al. (Pearce et al., 2022) prompted Copilot to generate code in scenarios related to high-risk security weaknesses (CWEs), covering multiple programming languages, and found that 40% of the generated code contained security vulnerabilities. Majdinasab et al. replicated the study by Pearce et al. using Python examples, and found that the newer versions of Copilot improved in terms of the number of security weaknesses in the generated code. However, they still provided vulnerable code suggestions (Majdinasab et al., 2024). Gupta et al. studied ChatGPT from the perspective of cybersecurity, and they found that AI tools such as ChatGPT can be used by both the defenders (secure code generation, vulnerability detection) and attackers (attack payload generation, malware code generation) (Gupta et al., 2023). Fu et al. analyzed AI-generated code from GitHub projects to identify security weaknesses in these code snippets (Fu et al., 2025). The study shows a high likelihood of security weaknesses in AI-generated code found in public projects.
@@
Ethical Concerns with AI-Generated Code. Existing studies have explored ethical concerns with AI-generated code. A recent survey by Stalnaker et al. with developers highlighted the ethical concerns in AI-generated code primarily regarding intellectual property concerns related to copyright violation (Stalnaker et al., 2024). This study shows that developers have mixed opinions about AI-generated code’s copyright and ownership attribution. Xu et al. investigated the ethical concerns and code quality issues with AI-generated code (Xu et al., 2025). These concerns include students’ unethical use of AI-generated code, code plagiarism, and questions of code provenance. The results emphasized the importance of distinguishing AI-generated code from human-written code for code review and quality assurance processes.
@@
Comparative Summary. Existing studies did not study whether developers acknowledge AI-generated code, how they acknowledge it, and what motivates them to acknowledge the use of AI-generated code. We investigate the practices of self-declaring AI-generated code among developers, examine the ways developers use to self-declare AI-generated code, including how they write the self-declaration comments and how they manage its scope, and explore the reasons behind self-declaring AI-generated code. Our study also used a mixed-methods approach to study these practices. Compared to existing work, our study addresses a unique topic with a more comprehensive research methodology. We detail our data collection and research methodology below.
@@ 3. Research Methodology
Our study employs a mixed-methods approach comprising two main phases: (1) an analysis of self-declaration comments from AI-generated code collected from open-sourced projects, and (2) an industrial survey with practitioners regarding the practices of self-declaring AI-generated code. We first collected data about AI-generated code from real-world projects by mining GitHub repositories to answer our research question of how developers self-declare AI-generated code. We followed that with a practitioner survey to validate the results from the mining study while addressing the question of why developers self-declare AI-generated code. This methodological triangulation strengthens the validity of our findings by complementing the mining study results with developers’ perspectives from the industrial survey. Figure 2 shows an overview of our research methodology.
@@ 3.1. Research Goal and Questions
Motivation. Developers may include self-declaration comments in their projects to indicate AI-generated portions of code. These comments may contain important information about the generated code (for example, the tool and prompt used to generate the code). To answer this RQ, we mine GitHub repositories to locate code comments that may indicate if the code is AI-generated. We refer to those comments as self-declaration comments. RQ1 is further divided into five sub-RQs.
@@
This RQ focuses on studying the content of the self-declaration comments collected from the projects. Developers may share useful information about AI-generated code in the added code comments (e.g., the AI tool/model used, prompt to generate the code, additional code explanation, and code quality indications). The information shared in these comments may provide insights into how developers write self-declaration comments.
@@
This RQ focuses on analyzing the scopes of the self-declaration comments based on information such as where they are mentioned (i.e., the position of the comment) in the file. The answer to this RQ helps in understanding the common ways developers use to manage the scope of the self-declaration comments, which is valuable for accurately identifying AI-generated code.
@@
This RQ focuses on analyzing the frequency of self-declared AI-generated code. We analyze two key aspects: (1) the distribution of the self-declaration comments within individual source files and (2) the number of files containing self-declaration comments within a single repository. This information helps quantify the presence of self-declared AI-generated code within repositories, offering meaningful information on its adoption by developers in practice.
@@
Motivation. AI code generation tools can speed up development and boost developers’ productivity (Ebert and Louridas, 2023; Liang et al., 2024). Developers increasingly use AI code generation tools in practice (SO-, [n. d.]). Here, we aim to investigate whether developers self-declare any AI-generated code used in their projects. We also explore the underlying reasons for developers to self-declare and to not self-declare AI-generated code. The answer to this RQ helps understand the current situation of self-declaration of AI-generated code as a practice in software development. RQ2 is further divided into the following two sub-RQs:
@@
This RQ investigates how widely developers self-declare (or not) AI-generated code in their projects.
@@
This RQ explores the underlying reasons why developers self-declare (or not) AI-generated code in their projects.
@@ 3.2. Mining Study Design
To answer RQ1, we collected self-declared AI-generated code from GitHub repositories. We defined a self-declared AI-generated code snippet as one in which the developer explicitly acknowledges that the code was written with the assistance of an AI code generation tool. This self-declaration occurs in the form of code comments added anywhere in the code file. We collected self-declared AI-generated code from GitHub because it is the most widely used platform by developers to host code (Oct, [n. d.]). This allowed us to access code repositories and study projects written in various languages.
@@
We used a keyword-based search to locate code snippets generated by four well-known AI code generation tools: Copilot, ChatGPT, CodeWhisperer, and Code Llama. We prioritized Copilot and ChatGPT due to their popularity and widespread use by developers in practice (Liang et al., 2024; SO-, [n. d.]).
3.2.1. Data Collection
@@
The first step of the data collection process is to identify the search terms to retrieve AI-generated code from GitHub projects. For this purpose, we performed a pilot data search to decide on the most suitable search terms. We used the names of popular AI code generation tools to locate AI-generated code snippets. However, using only the tool names mainly returned the projects where developers shared their experience using AI code generation tools, and did not yield code snippets generated by these tools. We also examined the use of more generic search terms without specifying the name of the tools used (e.g., “AI-generated code”, “AI-assisted code”), which returned only a few AI-generated code snippets. We finally noted that using specific search terms that include the name of the tools (e.g., “generated by GitHub Copilot”, “written by ChatGPT”) can return more relevant results than using those generic search terms. The search terms appreciating AI code generation tools, such as “thanks Copilot” and “thanks ChatGPT” also returned relevant results.
@@
We performed the pilot search using GitHub’s REST API (cod, 2022). As a result, we identified AI-generated code at the repository and code label. The repository and code labels can be referred to as search filters provided by GitHub for easy code searching. We classified the results as being Repository label when the repositories have an explicit acknowledgment of using AI-generated code in the project description (e.g., in the README files). On the other hand, Code label results returned the code files and snippets containing the self-declaration comments about the AI-generated code. The first author classified each code snippet (e.g., whether it is AI-generated), and then the results were cross-validated with two other co-authors. Through this process, we noticed that the search results from the repository label were more likely to contain dummy and student practice projects (i.e., small experimental projects with little content). Considering this, we decided to limit our data collection to the search results from the code label to ensure the quality of our dataset.
@@
In the search results from the code label, we noticed that developers included self-declaration comments in different positions of the code files, acknowledging that the code is AI-generated. The first author collected the code snippets with self-declaration comments from search results that mention the self-declaration of AI-generated code.
@@
# Search Term Results 1 written by ChatGPT 132 2 generated by ChatGPT 111 3 written by Copilot 86 4 thanks ChatGPT 66 5 generated by GitHub Copilot 66 6 thanks Copilot 41 7 by GitHub Copilot 21 8 AI-generated code 19 9 solution by ChatGPT 13 10 from ChatGPT 11 11 help of GitHub Copilot 10 12 solved by GitHub Copilot 6 13 coded by GitHub Copilot 6 14 solution by GitHub Copilot 4 15 from GitHub Copilot 4 16 generated by CodeWhisperer 4 17 from AWS CodeWhisperer 1 18 created by GitHub Copilot 1 19 using GitHub Copilot 1
@@
Based on the results of the pilot search, we selected a list of 23 search terms. Of those search terms, 22 included the names of the AI code generation tools and phrases specific to code generation. Only one search term was generic and did not include the name of any specific AI code generation tool. The list of search terms is provided in Table 1, which also shows the number of AI-generated code snippets included in our dataset.

### v2

@@ Abstract.
AI code generation tools have gained significant popularity among developers, who use them to assist in software development due to their capability to generate code. Existing studies mainly explored the quality, e.g., correctness and security, of AI-generated code, while in real-world software development, the prerequisite is to distinguish AI-generated code from human-written code, which emphasizes the need to explicitly declare AI-generated code by developers. To this end, this study intends to understand the ways developers use to self-declare AI-generated code and explore the reasons why developers choose to self-declare or not. We conducted a mixed-methods study consisting of two phases. In the first phase, we mined GitHub repositories and collected 613 instances of AI-generated code snippets. In the second phase, we conducted a follow-up practitioners’ survey, which received 111 valid responses. Our research revealed the practices followed by developers to self-declare AI-generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. The reasons for self-declaring AI-generated code include the need to track and monitor the code for future review and debugging, and ethical considerations. The reasons for not self-declaring AI-generated code include extensive modifications to AI-generated code and the developers’ perception that self-declaration is an unnecessary activity. We finally provided guidelines for practitioners to self-declare AI-generated code, addressing ethical and code quality concerns.
@@ 1. Introduction
Artificial Intelligence (AI) techniques, especially those employing Machine Learning (ML) models and Large Language Models (LLMs), are revolutionizing software engineering practices. There has been increased adoption of these models to develop more accurate, faster, and scalable tools that assist software engineers in their development tasks (Yellin, 2023; Hou et al., 2024) (e.g., AI-assisted coding (Liang et al., 2024)). One area that has been largely transformed with the advancement in LLMs is automatic code generation (Chen et al., 2021; Hou et al., 2024). Developers increasingly rely on AI code generation tools in their written code (AIt, 2024; AIR, 2023). The most recent Stack Overflow Developer Survey (2024) shows that there has been a significant adoption of code generation tools in practice (SO-, 2024, 2025). This is partially due to the widely available coding-specific tools (e.g., GitHub Copilot (Cop, 2024) and CodeWhisperer (Cod, 2024)) and general-purpose LLMs (e.g., ChatGPT (Cha, 2024) and Gemini (Gem, 2024)) that can generate a large amount of code with little human intervention, faster with high efficiency (Dakhel et al., 2023; Tian et al., 2023).
@@
Figure 1. An example of self-declaration comments of AI-generated code (urban-fingerprinting)
@@
The main contributions of this work are: (1) a curated dataset of 613 self-declared AI-generated code snippets from open-source repositories on GitHub and performing a comprehensive analysis with a focus on self-declaration comments to understand the practices of self-declaring AI-generated code, (2) a practitioners’ survey to validate the findings of the mining study and explore the potential reasons behind self-declaring (or not) AI-generated code, and (3) guidelines for practitioners to self-declare AI-generated code based on the results of the mining study and the follow-up practitioners’ survey.
Paper Organization: Section 2 discusses related work to this study. Section 3 presents our research questions and outlines the methodology employed for our data collection (from the mining study and the practitioners’ survey). We present our results in Section 4, followed by the discussion of the findings and implications in Section 5. We then present potential threats to the validity in Section 6 followed by the conclusion in Section 7.
@@ 2. Related Work
Quality, Correctness and Security of AI-Generated Code. Several recent studies have also investigated the quality and correctness of AI-generated code. Liu et al. evaluated ChatGPT-generated solutions to programming problems and studied their correctness and quality (through static analysis) (Liu et al., 2024a). The results showed that ChatGPT could generate correct solutions almost 70% of the time but suffers from maintainability issues. Ouyang et al. studied ChatGPT code generation capabilities and found that the non-determinism of ChatGPT affects the correctness and consistency of the generated code (Ouyang et al., 2024). Clark et al. evaluated ChatGPT-generated code quality and found that ChatGPT can generate good quality code through iterative prompts, but the generated code was found to contain bugs (Tambon et al., 2025), citing that ChatGPT-generated code would usually require review before integration into projects. These results show that AI-generated code will not hinder the software development process. Bui et al. introduced a framework (called OPENIA) that uses the internal states of code-generation LLMs to evaluate the correctness of AI-generated code (Bui et al., 2025). This approach was found to outperform existing methods, yielding significant improvements in performance metrics and enabling more effective quality assurance in AI code generation. Yu et al. investigated LLM-generated code on GitHub, analyzing the characteristics of such code, its associated projects, and the nature of subsequent modifications (Yu et al., 2024). Their findings revealed that ChatGPT and Copilot dominate LLM-generated code on GitHub, that this code is generally short and low in complexity, and that it undergoes relatively few bug-related changes.
@@
Similarly, the security of the generated code is another aspect that has been widely researched. Pearce et al. (Pearce et al., 2022) prompted Copilot to generate code in scenarios related to high-risk security weaknesses (CWEs), covering multiple programming languages, and found that 40% of the generated code contained security vulnerabilities. Majdinasab et al. replicated the study by Pearce et al. using Python examples, and found that the newer versions of Copilot improved in terms of the number of security weaknesses in the generated code. However, they still provided vulnerable code suggestions (Majdinasab et al., 2024). Gupta et al. studied ChatGPT from the perspective of cybersecurity, and they found that AI tools such as ChatGPT can be used by both the defenders (secure code generation, vulnerability detection) and attackers (attack payload generation, malware code generation) (Gupta et al., 2023). Fu et al. analyzed AI-generated code from GitHub projects to identify security weaknesses in these code snippets (Fu et al., 2025). The study shows a high likelihood of security weaknesses in AI-generated code found in public projects.
@@
Ethical Concerns with AI-Generated Code. Existing studies have explored ethical concerns with AI-generated code. A recent survey by Stalnaker et al. with developers highlighted the ethical concerns in AI-generated code, primarily regarding intellectual property concerns related to copyright violation (Stalnaker et al., 2024). The study reveals that developers hold mixed opinions regarding the copyright and ownership attribution of AI-generated code. Xu et al. investigated the ethical concerns and code quality issues with AI-generated code (Xu et al., 2025). These concerns include students’ unethical use of AI-generated code, code plagiarism, and questions of code provenance. The results emphasized the importance of distinguishing AI-generated code from human-written code for code review and quality assurance processes.
Automatic Detection of AI-Generated Code. There has also been a body of work that explored automated approaches to detecting AI-generated code. Gurioli et al. proposed a multilingual code stylometry approach based on transformer models that can distinguish AI-generated code from human-written code across ten programming languages with high accuracy (Gurioli et al., 2025). Similarly, Li et al. introduced a feature-based method for identifying ChatGPT-generated code, supported by dataset cleansing and code transformation techniques to improve detection reliability (Li et al., 2023). In contrast, our study does not focus on the automatic detection of AI-generated code but rather on developers’ self-declaration of any AI-generated code used in their projects.
@@
Comparative Summary. Existing studies have not investigated whether developers acknowledge AI-generated code, how they acknowledge it, or what motivates them to acknowledge the use of AI-generated code. We investigate the practices of self-declaring AI-generated code among developers, examine the ways developers use to self-declare AI-generated code, including how they write the self-declaration comments and how they manage its scope, and explore the reasons behind self-declaring AI-generated code. Existing studies, such as the work by Yu et al., performed a large-scale empirical analysis of GitHub repositories (Yu et al., 2024), and the study by Stalnaker et al., conducted a large-scale survey and interviews with developers (Stalnaker et al., 2024). In contrast to the methodologies used in these existing studies, our study used a mixed-methods approach comprising GitHub repositories mining and the follow-up practitioners’ survey to gather evidence about developers’ practices when self-declaring AI-generated code. In comparison to existing work, our study addresses a unique topic using a more comprehensive research methodology. We detail our data collection and research methodology below.
@@ 3. Research Methodology
Our study employs a mixed-methods approach comprising two main phases: (1) an analysis of self-declaration comments from AI-generated code collected from open-sourced projects, and (2) a practitioners’ survey with practitioners regarding the practices of self-declaring AI-generated code. We first collected data about AI-generated code from real-world projects by mining GitHub repositories to answer our research question of how (RQ1) developers self-declare AI-generated code. We followed that with a practitioner survey to validate the results from the mining study (RQ1) while addressing the research question of why (RQ2) developers self-declare AI-generated code. This methodological triangulation strengthens the validity of our findings by complementing the mining study results with developers’ perspectives from the practitioners’ survey. Figure 2 shows an overview of our research methodology.
@@ 3.1. Research Goal and Questions
Motivation. Developers may include self-declaration comments in their projects to indicate AI-generated portions of code. These comments may contain important information about the generated code (for example, the tool and prompt used to generate the code). To answer this RQ, we mine GitHub repositories to locate code comments that may indicate if the code is AI-generated. We refer to those comments as self-declaration comments. We also confirmed the findings we obtained from the mining study (RQ1) with a follow-up survey with practitioners. RQ1 is further divided into five sub-RQs.
@@
This RQ focuses on studying the content of the self-declaration comments collected from the projects. Developers may share useful information about AI-generated code in the added code comments (e.g., the AI tool/model used, the prompt used to generate the code, additional code explanations, and code quality indications). The information shared in these comments may provide insights into how developers write self-declaration comments.
@@
This RQ focuses on analyzing the scope of the self-declaration comments based on information such as where they are mentioned (i.e., the position of the comment) in the file. The scope of the self-declaration comments impacts the traceability and maintainability of the AI-generated code. A broad scope (file-level) may compromise accurate traceability, while a granular scope (snippet-level) may lead to redundancy and inefficiency. By analyzing common developer practices, this RQ provides guidelines for effective scope management in self-declaring AI-generated code.
@@
This RQ focuses on analyzing two key aspects: (1) the distribution of the self-declaration comments within individual source files and (2) the number of files containing self-declaration comments within a single repository. This information helps quantify the presence of self-declared AI-generated code within repositories. The presence and frequency of self-declaration comments in GitHub repositories can serve as indicators of how widely developers have adopted the self-declaration practices for AI-generated code. This RQ aims to illustrate the trend in adopting these self-declaration practices over the years.
@@
RQ2: What proportion of developers self-declare AI-generated code in practice? Why or why not?
Motivation. AI code generation tools can speed up development and boost developers’ productivity (Ebert and Louridas, 2023; Liang et al., 2024). Developers increasingly use AI code generation tools in practice (SO-, 2024). Therefore, we aim to investigate what proportion of developers self-declare any AI-generated code used in their projects and how frequently they self-declare such code. We also explore the underlying reasons for developers to self-declare and not to self-declare AI-generated code. The answer to this RQ helps understand the current situation of self-declaration of AI-generated code as a practice in software development. RQ2 is further divided into the following two sub-RQs:
@@
This RQ investigates how widely (the proportion of) developers self-declare (or not) AI-generated code in their projects from our practitioners’ survey. The answer to this RQ aims to provide insights into the current state of self-declaration practices of AI-generated code among developers.
@@
This RQ explores the underlying reasons why developers self-declare (or not) AI-generated code in their projects. This RQ aims to identify the potential motivation of developers to self-declare AI-generated code and any barriers that lead to the decision of not self-declaring AI-generated code.
@@ 3.2. Mining Study Design
To answer RQ1, we collected self-declared AI-generated code from GitHub repositories. We defined a self-declared AI-generated code snippet as one in which the developer explicitly acknowledges that the code was written with the assistance of an AI code generation tool. This self-declaration is expressed in the form of code comments added anywhere in the file. We collected self-declared AI-generated code from GitHub because it is the most widely used platform by developers to host code (Oct, 2024). This allowed us to access code repositories and study projects written in various languages.
@@
We used a keyword-based search to locate code snippets generated by four well-known AI code generation tools: Copilot, ChatGPT, CodeWhisperer, and Code Llama. We prioritized Copilot and ChatGPT due to their popularity and widespread use by developers in practice (Liang et al., 2024; SO-, 2024; Yu et al., 2024).
3.2.1. Keyword Identification
@@
The first step of the data collection process is to identify the search terms to retrieve AI-generated code from GitHub projects. For this purpose, we performed a pilot data search to determine the most suitable search terms. We used the names of popular AI code generation tools to locate AI-generated code snippets. However, using only the tool names mainly returned the projects where developers shared their experience using AI code generation tools, and did not yield code snippets generated by these tools. We also examined the use of more generic search terms without specifying the name of the tools used (e.g., “AI-generated code”, “AI-assisted code”), which returned only a few AI-generated code snippets. We finally noted that using specific search terms that include the name of the tools (e.g., “generated by GitHub Copilot”, “written by ChatGPT”) can return more relevant results than using those generic search terms. The search terms appreciating AI code generation tools, such as “thanks Copilot” and “thanks ChatGPT” also returned relevant results. We also searched using warning search terms, such as “warning ChatGPT” and “caution Copilot”. However, we were not able to find a considerable number of results using those search terms. The four search strategies and the pilot search terms are listed in Table 1.
# Search Strategy Example Terms
1 Tool Names Only ChatGPT, GitHub Copilot, Code Llama, Amazon CodeWhisperer
2 Generic AI Terms AI-generated code, AI-assisted code, code generated by AI, code written with the help of AI
3 Tool-Specific Patterns generated by {tool name}, written by {tool name}, from {tool name}, by {tool name}, solution by {tool name}, help of {tool name}, coded by {tool name}
4 Warning and Thanking Terms thanks {tool name}, credit to {tool name}, caution {tool name}, warning {tool name}
Table 1. Search strategies and initial pilot search terms to identify the search terms for data collection
@@ 1 Tool Names Only ChatGPT, GitHub Copilot, Code Llama, Amazon CodeWhisperer
We performed the pilot search using GitHub’s REST API (cod, 2022). As a result, we identified AI-generated code at the repository and code label. We classified the results as Repository label when the repositories explicitly acknowledge using AI-generated code in their project descriptions (e.g., in the README files). On the other hand, we classified results as Code label when the code files and snippets contain self-declaration comments about the AI-generated code. The first author classified each code snippet (e.g., whether it is AI-generated based on self-declaration comments), and then the results were cross-validated with two other co-authors. Through this process, we noticed that the search results from the repository label were more likely to contain dummy and student practice projects (i.e., small experimental projects with little content). Considering this, we decided to limit our data collection to the search results from the code label to ensure the quality of our dataset.
@@
In the search results from the code label, we noticed that developers included self-declaration comments in different positions of the code files, acknowledging that the code is AI-generated. The first author collected the code snippets from search results that mention the self-declaration of AI-generated code.
In the pilot search phase, we collected a significant number of self-declared AI-generated code using search terms that included the names of the AI code generation tools and phrases specific to code generation. Therefore, out of the total 23 search terms finally selected for our data collection, 20 were based on Strategy 3 (see Table 1). The list included only 2 “thanking” terms from Strategy 4 because we also found relevant results using these two terms during the pilot phase. We only included one general search term without the AI tool name (Strategy 2) because we found limited results for such search terms. We could not find any results using the terms from Strategy 1, and we did not include any search terms from this strategy. Based on the results of the pilot search, 22 search terms included the names of the AI code generation tools. 17 out of the 22 search terms included names of ChatGPT and GitHub Copilot, and only 5 search terms included names of other tools.. The list of search terms is provided in Table 2, which also shows the number of AI-generated code snippets included in our dataset.
@@
# Search Term Results # Search Term Results 1 written by ChatGPT 132 13 coded by GitHub Copilot 6 2 generated by ChatGPT 111 14 solution by GitHub Copilot 4 3 written by Copilot 86 15 from GitHub Copilot 4 4 thanks ChatGPT 66 16 generated by CodeWhisperer 4 5 generated by GitHub Copilot 66 17 from AWS CodeWhisperer 1 6 thanks Copilot 41 18 created by GitHub Copilot 1 7 by GitHub Copilot 21 19 using GitHub Copilot 1 8 AI-generated code 19 20 by CodeWhisperer 0 9 solution by ChatGPT 13 21 by Llama 0 10 from ChatGPT 11 22 from Code Llama 0 11 help of GitHub Copilot 10 23 proposed by GitHub Copilot 0 12 solved by GitHub Copilot 6